A company manufatures an alarm clock.3 weeks ago the company had 250 on hand 2 weeks from now the company will have 500.Assume the company will continue to make the clocks at the same rate. To writ the expresson,define a variable for the time in weeks from today and use this variable to writ a rule for the number of clocks manufactured.I have used so many different ways to write this and even went threw a math tutor site that gave me an answer that it wouldn't take.250=(50)(-3)+b 250=(50)(-3)+400 500=(50)2+400 we just can't figure out what it wants for the expression.
Answer by stanbon(75887)

A company manufatures an alarm clock.3 weeks ago the company had 250 on hand 2 weeks from now the company will have 500.Assume the company will continue to make the clocks at the same rate. To writ the expresson,define a variable for the time in weeks from today and use this variable to writ a rule for the number of clocks manufactured.I have used so many different ways to write this and even went threw a math tutor site that gave me an answer that it wouldn't take.250=(50)(-3)+b 250=(50)(-3)+400 500=(50)2+400 we just can't figure out what it wants for the expression.
--------------------
You have two points relating week and # of clocks: (-3,250) and (2,500)
With these you can determine a linear equation:
slope = [500-250]/[2--3] = 250/5 = 50
---------
Equation form is # of clocks = slope*week + b
Substitute one of the point values and the slope value to solve for "b":
250= 50*-3+b
b = 250+150 = 400
EQUATION:
# of clocks = 50(week) + 400
==============
You had the answer; you just didn't know you had the answer.
===============
Cheers,
Stan H.